Two automobiles start out at the same time from cities 595 km apart.if the rate of one is 8/9 of the rate of the other and if they meet in 7 hours,what is the rate of each?
Answer by mananth(16949)

Car I rate = xmph
car II rate = 8x/9 mph
They are moving towrads each other so add up their speeds
x+(8x/9)= 17x/9
distance d = 595
rate = 17x/9
time = 7
d=rt
595= 7*17x/9
multiply equation by 9
595 *9 = 119x
x= 595*9/119
speed of Car I = 45 mph
speed of Car II = 8*45/9= 40 mph
m.ananth@hotmail.ca